OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (accessdenied):

If you solve for y, you will have it in slope-intercept form. This can be done through some basic algebra -- division, subtraction... Slope-intercept form: y = mx + b m is the slope, and b is the y-intercept.

OpenStudy (anonymous):

So what is the answer? :|

OpenStudy (accessdenied):

5x - 3y = -6 -3y = -6 - 5x y = (-6)/(-3) - ((5)/(-3))x y = 2 + (5/3)x y = (5/3)x + 2 Slope: 5/3 y-intercept: 2